A telephone Booth 210cm tall casts a shadow 600cm long.at the same time, a nearby fire hydrant casts a shadow 240cm long.find the height of the fire hydrant.
1
Expert's answer
2020-10-14T18:46:26-0400
1. We will use the triangle similarity theorem because 2 triagles has equal angle between the hypotenuse of the triangle and its lower leg. This angle equal to the height of the sun.
2. By this theorem: height of telephone Booth/its shadow long = 
height of fire hidrant/its shadow long 
3. 210/600 = height of hidrant/240
4. h  = 240*210/600 =  50400/600  = 84 cm